OpenStudy (anonymous):

PLEASE HELP!!!!! Input in standard form the equation of the given line. The line that passes through (1, 1) and (3, 4)

OpenStudy (anonymous):

Ok, you need to know that the standard for is \[y=mx+c\]where m is the gradient of the line and c the y-intercept. There's a formula to find the gradient: \[m=\frac{ y _{2}-y_{1}}{ x _{2}-x_{1} }\] After you've found the gradient, you simply need to pick a point, say (1,1) and input in the standard form equation the values you have for y, x and m and solve the equation to find c.

OpenStudy (anonymous):

could you show the work to me?

OpenStudy (anonymous):

No, but I can show you the work for a made up example (: Find the equation of the line going through the points (1,2) and (2,6). Okay so we know we're dealing with a straight line, so the equation will have the form y=mx+c Now let's find m: we know the equation m=(y2-y1)/(x2-x1) So let's input our values: m=(6-2)/(2-1) m=4/1 m=4 Great we found m! Now let's find c. We know the line goes through the point (1,2) so the equation must hold true when x=1 and y=2; We know y=mx+c From what I just said we can write 2=1m+c We found m earlier 2=4+c c=2-4 c=-2 The equation of the line is therefore y=4x-2

OpenStudy (mathmale):

The easiest formula for the equation of a straight line, when the slope is known and one or two of the points is/are known, is \[y-y _{0}=m(x-x _{0}).\] I'd suggest that you do this first. Then, re-write this equation in the form y=mx + c, as Quantal has suggested.

OpenStudy (anonymous):

OK,thanks guys.this really helped:)